OpenStudy (anonymous):

Angel Falls is 750 meters higher than Niagara Falls. If both were 7 meters lower, than Angel Falls would be 16 times the height of Niagara. How high are each of them?

OpenStudy (gorv):

Angel =niagara+750-------(1) if height of both is reduced by 7 then difference in height will remain same but angel-7=16*(niagara -7) angel= 16*(niagara -7) + 7 ----(2) put 2 in 1 16*(niagara -7) + 7=niagara+750 16*niagara-112+7=750+niagara 15*niagara=750+105=855 niagara=855/15=57

OpenStudy (gorv):

angel=57+750=807

OpenStudy (gorv):

@darkknight321

OpenStudy (anonymous):

ok thank !